# Emergence of Abstractions: Concept Encoding and Decoding Mechanism for In-Context Learning in Transformers

### v2

@@ 3.4 Concept Encoding-Decoding Mechanism
Based on the observations above, we introduce concept encoding-decoding mechanism that serves as the core theory throughout the paper. The formal definition is in Appendix B
Definition 1 (Concept Encoding-Decoding Mechanism)
@@
Over the course of training, transformers learn separable representations by concept – concept encoding. Simultaneously, the model learns and applies concept-specific algorithms by leveraging the separable representation spaces – concept decoding. We define concept encoding-decoding mechanism as this two-stage process for ICL, which is mutually dependent with one another and emerges concurrently during training.
@@ 4 Towards Natural Experiments
In this section, we empirically validate the proposed concept encoding-decoding mechanism in pretrained LLMs. Specifically, we test several hypotheses derived from our proposed mechanism: whether pretrained LLMs exhibit concept encoding-decoding behavior and whether the quality of concept encoding-decoding predicts ICL performance on more natural tasks.
@@
We construct two classes of algorithmic tasks – natural language processing and arithmetic – comprising a total of 12 tasks. Within each class, the tasks are designed to be semantically similar, ensuring that the input distributions are alike across tasks. While the underlying latent concepts differ (e.g., different arithmetic operations or linguistic patterns), the surface features of the inputs remain consistent. By keeping the input distributions similar, we can effectively assess the model’s ability to infer and encode latent concepts based solely on subtle differences in the data, rather than the input variations. Refer to Appendix E for more details.
Part-of-Speech (POS) tagging. We construct a POS tagging (Blevins et al., 2022; Banko & Moore, 2004) dataset from Marcus et al. (1994), consisting of POS tags, such as Noun, Adjective, Verb, Adverb, Preposition, Pronoun. Given an input text and hidden POS tag $z_{i}$ (e.g., Noun), one needs to output the first word that is of the specified POS tag.
@@
We evaluate the existence of our proposed concept encoding-decoding mechanism across models of different families and scales (Gemma-2 2B/9B/27B and Llama-3.1-8B/70B) in Section 4.1 and Appendix E. We continue further analysis with the pretrained Llama-3.1-8B model (Meta, 2024). We do not train this model, except when we study the causal effect of concept decodability by finetuning in Section 4.3. We further detail the experimental setup in Appendix E.
@@
Concept Decodability (CD).
To quantify how well latent concepts can be inferred from representations, we employ a simple $k$-Nearest Neighbor (k-NN) classification metric. Inspired by prior studies using linear probes (Rimanic et al., 2020; Alain & Bengio, 2018), we assess whether the latent concepts can be extracted in a simple manner from their representations. Specifically, we use the representations of the token immediately before $y_{\ast}$ at a chosen layer and predict the latent concept by majority voting among its $k$ nearest neighbors ($k=10$, $N=100$).
@@
Figure 4: Concept encoding in Llama-3.1-8B. UMAP of the intermediate representations at layers 15 and 13 respectively for POS tagging and bitwise arithmetic with varying number of in-context examples (1, 4, 10-shot). With more in-context examples, the model builds increasingly separable representations clustered by their latent concepts.
4.1 Concept Encoding-Decoding in Pretrained LLMs
Hypothesis: Concept encoding-decoding behavior exists in pretrained LLMs.
We investigate the above hypothesis in two steps: (1) We first examine whether concept encoding occurs in pretrained LLMs; (2) We conduct mechanistic intervention studies to verify that different concept encoding triggers separate decoding algorithms, completing the full study of the concept encoding-decoding mechanism.
Step 1: Concept Encoding.
We first study whether the concept encoding occurs in pretrained LLMs. We vary the number of in-context examples for the different tasks and visualize the intermediate representations at the middle layers with UMAP in Figure 4. Given only 1-shot, where the model is expected to be confused about the latent concept, all of the representations are clustered and overlap with the Null class, which has no task latent. As examples increase, clustering by latent concepts emerges, becoming clearer by 10-shots. Interestingly, the separation of concepts, such as AND, OR, Noun, and Pronoun, is more pronounced. On the other hand, XNOR and XOR in bitwise arithmetic and Adjective and Preposition in POS tagging overlap significantly with Null. We conjecture the model likely sees and learns the former set of concepts better during pretraining. Overall, the analyses highlight that pretrained LLMs also exhibit concept encoding behavior and suggest taht in-context examples materialize as learning signal by creating more separable representations per latent concept . We will further explore this connection between the separability of concepts in the representation space and ICL performance in Section 4.2.
To quantify how the separability of representations translates into the decodability of the latent concepts, we compute the CD scores across the layers. In Figure 5(a), we see that the the decodability of latent concepts peaks in the intermediate layers, suggesting that the models indeed encode latent concepts through separable representations.
(a) CD Scores By Layers
(b) POS Tagging
(c) Bitwise Arithmetic
Figure 5: CD Scores by layers and number of demonstrations. (a) Mean CD scores across layers for POS tagging and Bitwise arithmetic with 4-shot in-context examples, showing peak decodability in intermediate layers. (b) For POS tagging and (c) for Bitwise arithmetic, CD scores all increase with the number of demonstrations, but the improvement in CD noticeably varies by task.
Step 2: Mechanistic Intervention Study.
Having shown concept encoding in pretrained LLMs, we now question whether encoded concepts trigger distinct decoding algorithms and whether they are causally related via mechanistic intervention studies Hendel et al. (2023); Todd et al. (2023). In this study, we hypothesize that, if the concept encoding is causally related to different decoding algorithms, helping or hindering the model’s ability to infer the latent concept in the given input should improve or degrade its performance in downstream tasks, respectively. We conduct this casual analysis by patching the output activations of a layer with the mean activations of 100 samples with the true latent concept (positive intervention) and with the Null latent variable (negative intervention). We present the results in Figure 18 of Appendix E.1. For POS tagging (Figure 18(a)), intervening positively improves performance by $\sim 14$% and intervening negatively degrades performance by $\sim 15$% on average across the 6 tasks. In bitwise arithmetic, the influence of interventions is less stark. Positive intervention improves performance by $\sim 2$% and negative intervention degrades performance by $\sim 6$% on average across all the 6 tasks. Both positive and negative interventions are more effective for tasks whose representation sub-spaces are clearly separated. For tasks whose representation overlap with those of Null, we hypothesize that the model is failing to infer the latent concept, rendering the intervention less effective.
Overall, through these two studies of the geometry of representations and mechanistic intervention, we demonstrate that concept encoding is causally linked to concept-specific decoding algorithms and that concept encoding-decoding behavior also exists in pretrained LLMs.
@@ 4.1 Concept Encoding-Decoding in Pretrained LLMs
Figure 6: CD score vs ICL performance in LLAMA-3.1 8B. We observe a positively correlated trend across most tasks. The grey dashed lines are linear lines of best fit. These results suggest that the accuracy of concept encoding is closely coupled with downstream ICL performance.
4.2 Predictability of In-context learning Task Performance
Hypothesis: Quality of concept encoding-decoding, measured by CD, is predictive of ICL performance.
We now investigate the second hypothesis of whether the quality of concept encoding-decoding is predictive of downstream ICL performance. If the model is conditionally applying a decoding algorithm to perform the task by first inferring the latent concept, the accuracy of the latent concept encoding (measured by CD) and ICL task performance should be closely correlated. To this end, we analyze the relationship between CD and test accuracy in Figure 6. In both datasets, we see that, generally, higher CD scores correspond to better performance on the respective tasks. More interestingly, referring back to Figure 4, we again remark that the representations of some classes (Adjective and Preposition in POS tagging and XOR and XNOR in bitwise arithmetic) are mapped to those of the Null class. We notice that this set of classes whose representations overlap with those of Null generally have low task performance and do not improve as much as the others given more demonstrations. We conjecture that the model does not accurately encode latent concepts of those that are overlapped with the Null class representations.
We also test the generality of the predictability of ICL performance from CD across a different model family (Team, 2024) and scales. We conduct the same analysis on Gemma-2 2B, 9B, and 27B and Llama-3.1 70B and present the results in Figure 19 in Appendix E.2. These results demonstrate that the correlation between CD and ICL performance is robust across models and tasks. Interestingly, in all of the Gemma-2 family and Llama-3.1 70B models, Noun, Pronoun, and Verb show the clearest signs of concept encoding-decoding behavior, as we saw in the Llama-3.1 8B model. In the bitwise arithmetic task, AND, NAND, OR, and NOR (classes that showed the strongest encoding-decoding behavior in Llama-3.1 8B), also show the strongest signs of concept encoding-decoding behavior across all of these models. Given that many LLMs are trained on similar sources of pretraining data (Soldaini et al., 2024; Gao et al., 2020) (CommonCrawl, Wikipedia, etc.), we conjecture that the models may have learned similar encoding-decoding mechanisms for these concepts.
Another natural curiosity that arises is whether this correlation can also be observed during pretraining. Although computationally infeasible to explore this by conducting large-scale pretraining, we demonstrate the correlation between CD and performance by evaluating them over the course of OLMo-7B pretraining (Groeneveld et al., 2024) in Figure 9 of Appendix A. Overall, these results demonstrate that the model’s ability to infer the correct latents is generally correlated to its ICL task performance.
(b) POS Tagging
Figure 7: (a) CD scores across layers for POS and arithmetic after finetuning the first 10 and last 10 layers, at 4 shots. While finetuning (FT) the last 10 layers has minimal effect on the CD scores, finetuning the first 10 layers significantly increases the concept decodability. This phenomenon is accompanied by noticeable improvement in ICL performance. PT denotes the pretrained LLM. (b) UMAP visualization of FT first 10 layers. We illustrate that the increased CD scores correspond to a clear cluster of the representations by latent concepts.
4.3 Investigating the Causal Effect of Concept Encoding by Finetuning
Hypothesis: In transformer-based LLMs, earlier layers learn to encode concept, whereas the latter layers condition the algorithm on it. Thus, finetuning only the earlier layers can improve concept encoding and thus will be more effective for improving ICL performance than finetuning only the latter layers.
@@ 4.3 Investigating the Causal Effect of Concept Encoding by Finetuning
To further investigate the causal importance of concept encoding for downstream ICL performance, we perform two types of finetuning: only the first 10 layers versus only the last 10 layers. We previously found that concept encoding occurs in the middle layers (layer 15 for POS tagging and layer 13 for bitwise arithmetic). Finetuning only the last 10 layers restricts the model from learning to encode latent concepts in intermediate representations. As illustrated in Figure 7, finetuning the last 10 layers barely changes their CD scores from the pretrained model. In contrast, finetuning the first 10 layers significantly improves the CD scores and aligns the representation subspaces with the inferred latent concepts. This improvement in CD scores directly translates to significantly better ICL task performance. With 4-shot examples, finetuning the first 10 layers outperforms finetuning the last 10 layers by $\sim$ 37% in the POS task and 24% in bitwise arithmetic. In the bitwise arithmetic task, finetuning the first 10 layers achieves near-perfect accuracy for all tasks except XNOR, whose representations overlap with those of Null.
(a) POS Tagging
(b) Bitwise Arithmetic
4.4 Investigating the Effect of Prompting on Concept Encoding-Decoding
@@
Hypothesis: Prompting enhance CD by providing a stronger learning signal for concept inference, and thus improve ICL performance correspondingly.
Our study reveals that enhancing concept encoding is a unifying principle that improves in-context learning (ICL) across different strategies. We observe in Sections 4.1, 4.2, and 4.3 that increasing in-context examples and finetuning facilitate building separable representations by their latent concepts. Prompting (Wei et al., 2023) is also a simple and common method to improve a model’s ICL performance. Thus, we experiment with prompting as part of our investigation. We question whether providing the underlying concept (i.e., including true labels of bitwise arithmetic) indeed enhances concept encoding and, as expected, performance. As shown in Figures 21 and 22 of Appendix F, prompting in fact improves the concept encoding and performance simultaneously. However, we interpret these results with caution, since the model may be capturing spurious correlations from the input prompt differences.

### v3

@@ 3.4 Task Encoding-Decoding
Based on the observations above, we define task encoding and decoding that serves as the core framework throughout the paper. The formal definition is in Appendix B.
Definition 3.1 (Task Encoding and Decoding).
@@
Over training, transformers learn separable representations by latent tasks – task encoding. Simultaneously, the model learns task-specific algorithms by leveraging the separable representation spaces – task decoding. We illustrate that these two processes combined manifest as task vectors, and that they emerge concurrently during training.
@@ 4 Towards Natural Experiments
In this section, we empirically study the task encoding-decoding phenomenon in pretrained LLMs. Specifically, we test several hypotheses driven by the the proposed task encoding-decoding framework and demonstrate that our geometric measure - called Task Decodability - can serve as a proxy of the quality of task vector and ICL performance.
@@
We construct two classes of algorithmic tasks – natural language processing and arithmetic – comprising a total of 12 tasks. Within each class, the tasks are designed to be semantically similar, ensuring that the input distributions are alike across tasks. While the underlying tasks differ (e.g., different arithmetic operations or linguistic patterns), the surface features of inputs remain consistent. By keeping the input distributions similar, we can effectively assess the model’s ability to infer and encode latent concepts based solely on subtle differences in the data, rather than simply relying on the input variations. Refer to Appendix G for more details.
Part-of-Speech (POS) tagging. We construct a POS tagging (Blevins et al., 2022; Banko & Moore, 2004) dataset from Marcus et al. (1994), consisting of POS tags, such as Noun, Adjective, Verb, Adverb, Preposition, and Pronoun. Given an input text and hidden POS tag $z_{i}$ (e.g., Noun), one needs to output the first word that is of the specified POS tag.
@@
We use Llama-3.1-8B as the default model for the experiments. We test our hypothesis across different families and scales (Gemma-2 2B/9B/27B and Llama-3.1-8B/70B) in Section 4.1 and Appendix G. For analyzing the models over the course of pretraining, we use the checkpoints of OLMo 7B (Groeneveld et al., 2024). We do not train any model, except when we study the causal effect of task decodability by finetuning in Section 4.4. We further detail the experimental setup in Appendix G.
@@
Task Decodability (TD).
To quantify how the task encoding progresses, we employ a simple $k$-Nearest Neighbor (k-NN) classification metric. Inspired by prior studies using linear probes (Rimanic et al., 2020; Alain & Bengio, 2018), we assess whether the latent task can be extracted in a simple manner from their representations. Specifically, we use the representations of the token immediately before $y_{\ast}$ at a chosen layer and predict the latent task by majority voting among its $k$ nearest neighbors ($k=10$, $N=100$). For choosing which layer to measure TD, we compute TD across all layers and choose the layer that best encodes the task. We empirically validate with UMAP visualizations that the representations become separable precisely when the TD score peaks (See Figures 10, 11 and 19).
We now formally define TD. Let $\mathcal{T}$ be the set of latent tasks. For each task $z\in\mathcal{T}$, sample $N=100$ datapoints $\{(x_{i},y_{i})\}_{i=1}^{N}$ and collect the intermediate representations $\{h_{i}\}_{i=1}^{N}$ from a chosen layer (e.g., the token embedding immediately after $x_{i}$). Label each representation $h_{i}$ with the corresponding task $z$. This yields a set $\mathcal{S}=\bigcup_{z\in\mathcal{T}}\{(h_{i},z)\mid i=1,\dots,N\}$ , over all tasks $z$.
Given a query point $(h_{i},z)$, we exclude $(h_{i},z)$ from $\mathcal{S}$ (i.e., $\mathcal{S}\setminus(h_{i},z)$) and find its $k$ nearest neighbors (with $k=10$) in the remaining set. We then use majority voting on these neighbors’ task labels to produce a predicted label $\hat{z}$. If $\hat{z}=z$, classification for the task label is correct. The TD score at this layer is the fraction of query points classified correctly. $\text{TD}\;=\;\frac{1}{|\mathcal{D}|}\sum_{(h_{i},z)\in\mathcal{D}}\mathbf{1}\bigl{[}\hat{z}=z\bigr{]}.$
Figure 3: Task encoding in Llama-3.1-8B. (Top) UMAP of the intermediate representations and (Bottom) TD scores for each tasks at layers 15 and 13, respectively, for POS tagging and bitwise arithmetic with varying numbers of in-context examples.
4.1 Task Encoding and Decoding in Pretrained LLMs
Hypothesis 1: In pretrained LLMs, task decodability varies across tasks and determines the effectiveness of the task vector.
Task decodability varies across tasks.
We begin by qualitatively analyzing the intermediate representations of LLMs across different tasks using UMAP. As shown in Figure 3, the degree of localization of representations depends on the task. For instance, tasks like AND, OR, Noun, and Pronoun exhibit clear and distinct clusters when sufficient in-context examples (e.g., 10-shots) are provided, while tasks like XNOR and XOR in bitwise arithmetic or Adjective and Preposition in POS tagging remain overlapped with the Null class. This suggests that the model inherently learns certain tasks better than others during pretraining, likely due to differences in their predictability or frequency in the pretraining corpus. Importantly, the emergence of separability with increasing examples highlights that in-context examples act as a signal for materializing task-specific representations.
To better quantify such variability of representational separations between different tasks in Figure 3, we use our proposed Task Decodability (TD) scores. Analogous to the UMAP visualizations, we observe that some tasks, such as Noun and Pronoun in POS tagging and AND and OR in bitwise arithmetic, are much more decodable from their representations than the others. Also, we confirm in Figure 19(a) of Appendix G.1, TD score peaks in the middle, suggesting earlier layers encode the task and latter layers execute the decoding algorithm.
Task decodability as an indicator of task vector effectiveness.
Through mechanistic interventions, we further examine how TD represents the quality of the task vector formation – how effective injecting the task vector is at triggering the desired ICL behavior. By patching layer outputs with either mean activations from correctly inferred latent concepts (positive intervention) or the Null class (negative intervention), we measure their impact on task performance. As shown in Figure 20, tasks with highly separable representations, such as Noun and Verb in POS tagging and AND and OR in bitwise arithmetic, exhibit significant improvements from positive interventions (up to $\sim 14\%$) and noticeable degradation from negative interventions (up to $\sim 15\%$).
In contrast, tasks with overlapping representations, such as XOR, XNOR, Adjective, and Preposition, show limited sensitivity to interventions, with positive interventions improving performance by only $\sim 2\%$ and negative interventions causing a modest $\sim 6\%$ drop. These differences suggest that when task encoding is done correctly and the representations are well-separated, the task vectors are more effective at guiding decoding processes, as they can easily map distinct concepts to their respective downstream algorithms. Conversely, overlapping representations hinder the ability of the task vectors to reliably trigger distinct decoding algorithms, reducing their overall effectiveness.
Overall, these findings demonstrate that task decodability is a direct reflection of task vector quality. This implies that, when the model becomes more adept at a task, the task becomes correspondingly easier to decode from their representations. We further assess this hypothesis in the next section.
@@ 4.1 Task Encoding and Decoding in Pretrained LLMs
Figure 4: TD score vs ICL performance in Llama-3.1 8B. We observe a positively correlated trend across most tasks. The grey dashed lines are linear lines of best fit. These results suggest that the accuracy of task encoding is closely coupled with downstream ICL performance.
4.2 Task Decodability Predicts ICL Task Performance
Hypothesis 2: Quality of task encoding-decoding is predictive of ICL performance.
We now investigate the second hypothesis of whether the quality of task encoding-decoding is predictive of downstream ICL performance. If the model is conditionally applying a decoding algorithm by first inferring the latent task, the quality of the latent task encoding (measured by TD score) and ICL task performance should be closely correlated. To this end, we analyze the relationship between TD and test accuracy in Figure 4. In both datasets, we see that, higher TD scores correspond to better performance on the respective tasks. Notably, referring back to Figure 3, we again remark that the representations of some classes (Adjective and Preposition in POS tagging and XOR and XNOR in bitwise arithmetic) are mapped to those of the Null class. We notice that this set of classes whose representations overlap with those of Null generally have low task performance and do not improve as much as the others given more demonstrations. We conjecture that the model does not accurately encode the latent tasks of those that are overlapped with the Null class representations.
We also test the generality of our hypothesis that TD predicts ICL performance across different model families and scales. We perform the same analysis on Gemma-2 2B, 9B and 27B (Google, 2024) and Llama-3.1 70B and present the results in Figure 21 in the Appendix G.3. These results demonstrate that the correlation between TD and ICL performance is robust across models and tasks. Interestingly, in all of the Gemma-2 family and Llama-3.1 70B models, Noun, Pronoun, and Verb show the clearest signs of task encoding and decoding behavior, as we saw in the Llama-3.1 8B model. In the bitwise arithmetic task, AND, NAND, OR, and NOR (classes that showed the strongest encoding-decoding behavior in Llama-3.1 8B), also show the strongest signs of task encoding-decoding behavior across all of these models. Given that many LLMs are trained on similar sources of pretraining data (Soldaini et al., 2024; Gao et al., 2020) (CommonCrawl, Wikipedia, etc.), we conjecture that the models may have learned similar task vector mechanisms for these concepts (Huh et al., 2024).
@@ 4.2 Task Decodability Predicts ICL Task Performance
Figure 5: TD scores and ICL accuracy during OLMo-7B pretraining (Groeneveld et al., 2024), averaged across POS tagging and bitwise arithmetic tasks, evaluated using 4-shot prompts on 1000 test examples per task. The shaded regions represent the standard deviation across the tasks.
4.3 Studying the Emergence of Task Encoding and Decoding in LLM Pretraining
Another natural question is whether the simultaneous emergence of ICL ability and task encoding observed in synthetic experiments in Section 3.3 also occurs during LLM pretraining. Since large-scale pretraining studies are computationally infeasible, we leverage different training checkpoints of OLMo-7B (Groeneveld et al., 2024) to investigate the relationship between TD and ICL task performance during pretraining. As shown in Figures 5 and 6, increases in TD scores closely align with gains in ICL accuracy. Interestingly, unlike the synthetic setup, this progression is more gradual. This suggests that task encoding and decoding for natural ICL tasks emerge more gradually during pretraining, likely because LLMs are simultaneously learning a diverse range of tasks, making the training dynamics more intricate to disentangle and comprehend. Thus, further investigation into how task-specific representations evolve during pretraining is warranted.
Generalization to Recurrent Language Models
Recent work shows that recurrent language models, such as LSTMs (Xie et al., 2021) or Mamba (Gu & Dao, 2023), exhibit ICL behavior (Grazzi et al., 2024). To verify whether our encoding-decoding view explains the emergence and effectiveness of task vectors in such models, we performed the same analysis on Mamba 8B (see Appendix G.4). As shown in Figure 22, a strong positive correlation between TD scores and task performance confirms our hypothesis that TD scores predict ICL performance for state-space models as well. This implies that the proposed task encoding-decoding perspective could potentially be generalized across various architectures beyond transformers.
4.4 Improving Task Encoding by Finetuning the Early Layers
Hypothesis 3: Building on the encoder-decoder framework, finetuning the early layers enhances task encoding and should yield greater improvements in ICL performance compared to finetuning the later layers.
POS Tagging Bitwise Arithmetic
Avg. TD Avg. Acc. Avg. TD Avg. Acc.
Pretrained 0.68 $\pm$ 0.11 25.5 $\pm$ 6.7 0.43 $\pm$ 0.13 26.2 $\pm$ 15.5
FT (first 10) 0.95 $\pm$ 0.02 69.4 $\pm$ 14.2 0.85 $\pm$ 0.11 90.2 $\pm$ 15.0
FT (last 10) 0.68 $\pm$ 0.11 33.9 $\pm$ 12.3 0.43 $\pm$ 0.13 66.3 $\pm$ 9.01
Prompting 0.97 $\pm$ 0.02 34.4 $\pm$ 10.3 0.90 $\pm$ 0.10 57.2 $\pm$ 30.1
Table 1: Average TD scores and task accuracies with finetuning and prompting for POS Tagging and Bitwise Arithmetic.
Given our observation that the earlier layers are responsible for task encoding and the latter layers for the decoding algorithms, we hypothesize that finetuning the earlier layers will enhance task representation and improve ICL task performance more effectively than finetuning the latter layers. This approach challenges the common assumption in model finetuning (Wu et al., 2024; Kumar et al., 2022), where the focus typically lies on adjusting the latter layers or the linear head, under the belief that the final layers primarily govern task-specific adaptation. However, our findings reveal that targeting the earlier layers provides a more effective pathway for improving ICL performance.
Finetuning only the last 10 layers results in minimal improvements to task encoding, as these layers do not change the upstream representations. As illustrated in Table 1, the TD scores of the last 10 layers remain unchanged compared to the pretrained model. In contrast, finetuning the first 10 layers significantly improves TD scores, aligning the representation subspaces with the latent concepts and enhancing task-specific abstractions.
@@ 4.4 Improving Task Encoding by Finetuning the Early Layers
This improvement in representation encoding directly translates to better downstream ICL performance. With 4-shot examples, finetuning the first 10 layers outperforms finetuning the last 10 layers by 37% in the POS task and 24% in bitwise arithmetic. Moreover, finetuning the first 10 layers achieves near-perfect accuracy across bitwise arithmetic tasks, except for XNOR, where overlapping representations with Null limit further improvement.
4.5 Investigating the Effect of Prompting on Task Encoding and Decoding
@@
Hypothesis 4: Prompting enhances TD by providing a stronger learning signal for task inference, and thus improves ICL performance correspondingly.
In previous sections, we showed that in-context examples and finetuning both improve task encoding and hence ICL task performance. Prompting (Liu et al., 2023) is also a simple and common method to improve a model’s performance. In this section, we experiment with prompting to see how it changes TD along with performance. As shown in Table 1, prompting in fact improves the task encoding and performance simultaneously. Together with the previous results, it suggests that enhancing task encoding may be the unifying principle through which the learning signal materializes in the model’s representations across different strategies (e.g., in-context examples, finetuning, and prompting). We detail the setup and results in Appendix I.